OpenStudy (anonymous):

rewrite in slope-intercept form 2x-3y=6

OpenStudy (anonymous):

slope intercept form is mx+b=y

OpenStudy (anonymous):

so solve for y by isolating it

OpenStudy (anonymous):

so it would go as follows 2x-3y=6 first subract the 2x to isolate the y 3y=6-2x then divide by 3 to further isolate the y y=2-2/3x